Circular No. 8179

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hc IN UGC 1993 Further to IAUC 8178, T. Boles reports the discovery of an apparent

supernova (mag 18.1) on unfiltered CCD images taken on Aug. 7.094 and 10.098 UT. The new object is located at α = 2^{h}31^{m}42^{s}.04, δ = +39^{o}22′05′′.5 (equinox 2000.0), which is ≈ 22′′.5 east and 36′′.6 south of the center of UGC 1993. SN 2003hc is not present on Boles’ images from Jan. 15, Feb. 5, and 21 (limiting mag 19.5), and it is not present on Palomar Sky Survey red or blue plates.

SUPERNOVA 2003hd IN MCG –04-5-10 Further to IAUC 8175, M. Ganeshalingam and W. Li report the LO-

TOSS discovery of an apparent supernova on unfiltered KAIT images taken on Aug. 9.5 (mag ∼ 17.4) and 10.5 UT (mag ∼ 17.3). SN 2003hd is located at α = 1^{h}49^{m}46^{s}.31, δ = −21^{o}54′37′′.8 (equinox 2000.0), which is 11′′.0 west and 10′′.2 south of the nucleus of MCG –04-5-10. A KAIT image taken on Jan. 1.2 showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2003 K5 AND C/2003 K6 (SOHO) Further to IAUC 8178, D. Hammer has reported astrometric measure-

ments for two additional Meyer-group comets found on C2 coronagraph images at the SOHO website by X. Leprette (C/2003 K5) and X.-M. Zhou (C/2003 K6). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2003-P02.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 K5 May 21.201  3^{h}53.^{m}4  +21^{o}06′
C/2003 K6  30.061  4 28.5  +23 10

COMET C/2003 F1 (LINEAR)

Improved orbital elements from MPC 49276:

Epoch = 2003 June 10.0 TT

T = 2003 June 28.4234 TT  ω = 121.^{o}1818
e = 0.806021  Ω = 87.4880 }2000.0
q = 4.007834 AU  i = 70.2214

a = 20.661125 AU n^{o} = 0.0104948 P = 93.91 years

2003 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green